How do i determine the equation of this line: parallel to the line -2x+3, containing the point(-2,1)?
Answer by Pieter(35)

parallel to -2x + 3 means: with the same slope! so the slope of your needed line is also -2.
You have to use y - y1 = m (x - x1) whete m = slope (-2 here), and x1 and y1 are a point(x1,y1) (here (-2,1)) where the line has to go through.
We get: y - 1 = -2 (x + 2)
or y - 1 = -2x - 4
or y = -2x - 3
This is the equation of the line you needed to determine!
Greets,
Pieter Gurdebeke